# Improving accuracy of tree-tensor network approach by optimization of network structure

### v1

@@ Abstract
Numerical methods based on tensor networks have been extensively explored in the research of quantum many-body systems in recent years. It has been recognized that the ability of tensor networks to describe a quantum many-body state crucially depends on the spatial structure of the network. In the previous work, we proposed an algorithm based on tree tensor networks (TTNs) that automatically optimizes the structure of TTN according to the spatial profile of entanglement in the state of interest. In this paper, we precisely analyze how detailed updating schemes in the structural optimization algorithm affect its computational accuracy for the random XY-exchange model under random magnetic fields and the Richardson model. We then find that for the random XY model, on the one hand, the algorithm achieves improved accuracy, and the stochastic algorithm, which selects the local network structure probabilistically, is notably effective. For the Richardson model, on the other hand, the resulting numerical accuracy subtly depends on the initial TTN and the updating schemes. In particular, the algorithm without the stochastic updating scheme certainly improves the accuracy, while the one with the stochastic updates results in poor accuracy due to the effect of randomizing the network structure at the early stage of the calculation. These results indicate that the algorithm successfully improves the accuracy of the numerical calculations for quantum many-body states, while it is essential to appropriately choose the updating scheme as well as the initial TTN structure, depending on the systems treated.
@@ I Introduction
Since the development of the density-matrix renormalization-group (DMRG) method[1, 2], tensor-network approach has made rapid progress in a variety of fields[3, 4, 5]. While tensor networks have been studied in statistical mechanics for long years[6], the introduction of quantum-information concepts has advanced the understanding of their properties significantly, and it has been recognized that tensor networks can represent the low-energy states of quantum many-body systems efficiently. Several algorithms based on various tensor-network states, including matrix-product state (MPS)[7, 8], tensor-product state (TPS)[9, 10] or projected entangled-pair state (PEPS)[11, 12], and the state of multiscale entanglement renormalization ansatz (MERA)[13, 14], have been developed.
@@
In this paper, we explore a numerical algorithm based on the tree-tensor networks (TTNs), which are tensor networks without loops. We particularly focus on binary trees, TTNs composed of three-leg tensors, although the extension to $m$-ary trees is straightforward. An essential property of the TTN is that it can be separated into two parts by cutting an arbitrary bond in the network. This property of TTN leads to the following guideline for reducing the loss of accuracy by truncation.[15, 16, 17] Since every bond in a TTN has a corresponding bipartition of the system, each bond must be able to represent the quantum entanglement between the subsystems of the corresponding bipartition. However, due to the upper bound of the bond dimension, which we denote by $\chi$ in the following, the entanglement entropy (EE) that a bond can carry is also upper bounded to a maximum of $\ln\chi$. If the EE associated with each bond exceeds the upper limit, the excess entanglement is omitted in the approximated TTN and results in a loss of accuracy. Here, recall that the number of bonds in a TTN, which is $32N-$ 3 where $N$ is the number of quantum spins in the system, is much smaller than the number of the possible bipartitions of the system, $12^{N-1}-$ 1. Therefore, the network structure that makes the EEs on the $(2N-3)$ bonds as small as possible can minimize the loss of accuracy in approximating the target quantum state by a TTN with a finite $\chi$. We call this guideline to minimize the EEs on a TTN the least-EE principle.
@@
The least-EE principle indicates that using a TTN with the optimal structure is crucially important for improving the accuracy of the TTN approach. However, finding the optimal TTN structure is highly nontrivial and challenging. Several efforts have been devoted to the problem so far. There have been attempts to optimize the order of sites within the matrix-product network (MPN), which is employed in DMRG, based on various guidelines.[18, 19, 20, 21, 22] A method to optimize the structure of TTN using the cut-off bond dimension as a minimization function has also been proposed.[23] For systems with randomness, the tensor-network strong-disorder renormalization group (tSDRG) method, in which a TTN suitable for a random system of interest is constructed based on the idea of the strong-disorder renormalization group,[24, 25] has been developed[26, 27] and successfully applied.[28, 29, 15] For small systems for which exact wavefunction is obtained, a method to extract the optimal TTN structure from the wavefunction according to the least-EE principle has also been proposed.[17]
@@
In the previous study[16], we proposed an algorithm to search for the optimal TTN structure by iterating the local reconstruction of the network based on the least-EE principle. We applied the algorithm to some toy models for which the optimal TTN structure was deduced from the perturbative renormalization analysis[16, 30]. We then demonstrated that the algorithm succeeded in obtaining the expected optimal structure. In this paper, we investigated the performance of several types of the TTN structural optimization calculations, which are specified by the initial TTN structure, the scheme to select the local TTN structure in the optimization process, and the bond dimension at which the structural optimization is executed. We performed the calculations of the types on the random XY-exchange model under random magnetic fields and the Richardson model. We thereby demonstrated that for the random XY model, the structural optimization algorithm could reduce the variational energy significantly. In particular, stochastic schemes, which select the local TTN structure probabilistically, achieved the lowest variational energy regardless of the initial TTN structure. For the Richardson model, we found that while the structural optimization algorithm could yield a good TTN structure with a low variational energy, the calculations were sometimes trapped by a TTN structure unsuitable for representing the lowest energy state and failed to reach the optimal TTN structure. These results for the two models reveal that while the TTN structural optimization algorithm is basically effective in improving the accuracy of the calculation, the performance of each type of the calculations depends on the model. The results thus indicate the importance of appropriately choosing the updating scheme of the network and the initial TTN structure according to the model to be studied.
@@
The plan of this paper is as follows. In Sec. II, we review the structural optimization algorithm and make several comments. Section III.1 describes the types of the optimization calculation tested in the present study and the quantities calculated. Sections III.2 and III.3 respectively present the numerical results for the random XY-exchange model under random magnetic fields and the Richardson model. The performances of each type for the models are discussed there. Section IV is devoted to a summary and concluding remarks. In Appendix A, we explain how to prepare the initial TTN. The details of the actual calculation, including the number of sweeps, are presented in Appendix B. Appendix C describes the exact calculation of the lowest energy for the Richardson model.
@@ II Algorithm
Figure 1: Examples of TTN. (a) Matrix-Product Network (MPN), (b) Perfect-Binary Tree (PBT), (c) TTN of a generic form, and (d) the same TTN as that in (c). Open circles represent isometries and the solid diamond represents the singular-value matrix. Circles with colors represent bare spins. In (a), (b), and (c), the bare spins are arranged in the order of the site index. Note that within the category of MPN and PBT, there are many networks with different site orderings.
Figure 2: (a) Isometry $V_{a,b}^{c}$ with the legs $a$, $b$, and $c$. (b) Schematic picture of the orthonormal condition Eq. (1). Connected legs (bonds) represent a contraction of their degrees of freedom. The connected arrows in the right-hand-side of (b) represents a Kronecker delta $\delta_{cc^{\prime}}$.
@@
where $\bar{V}_{a,b}^{c}$ is a complex conjugate of $V_{a,b}^{c}$, and call the tensor the isometry. We assign a direction to each bond in the TTN so that the bases of the bond outgoing from each tensor are orthonormal; Namely, in Eq. (1), the index $c$ ($c^{\prime}$) corresponds to the outgoing bond, while the indices $a$ and $b$ correspond to the incoming bonds. All bonds in the TTN are directed from the boundary, where the bare spin degrees of freedom are, towards the “center” of the TTN, where the diagonal matrix of singular values is located. Such a representation format of the TTN is called the mixed canonical form[31]. Adopting the mixed canonical form offers strong advantages, such as guaranteeing the orthonormality of the basis of the truncated Hilbert space expanded by incoming bonds and making the contraction caluculations of isometries extremely simple as shown in Fig. 2 (b). We note that the center position of the canonical form can be taken at an arbitrary bond and can also be moved by the fusing and re-decomposing process of tensors[31].
Figure 3: (a) Central area at the biginning of a step. (b) Effective ground-state wave function $\Psi(abcd)$ obtained by the diagonalization of the effective Hamiltonian $\tilde{\mathcal{H}}$. (c), (d), and (e) show the possible local connections of the singular-value decomposition Eqs. (2a), (2b), and (2c), respectively. Open circles represent isometries and a solid diamond denotes the singular-value matrix at the canonical center.
@@
2. Construct the effective Hamiltonian $\tilde{\mathcal{H}}$ of the whole system on the orthogonal basis of the central area.
@@
4. Perform the singular-value decomposition of $\Psi$ for the three possible local connections in Figs. 3 (c)-(e).
@@ 5. Choose the optimal local connection according to the adopted updating scheme of the network discribed
Throughout this paper, we distinguish the isometries by the letters for the bond index, if necessary. [Hence, for example, $V^{e}_{ab}$ and $V^{e}_{cd}$ in Eq. (2a) are different isometries.] We also adopt the notation that the singular values are arranged in the descending order, $D_{1}\geq D_{2}\geq\cdots\geq D_{\chi^{2}}\geq 0$ (the same for $\{D^{\prime}_{e}\}$ and $\{D^{\prime\prime}_{e}\}$) and suppose that the ground-state wave function $\Psi(abcd)$ is normalized so that $\sum_{e=1}^{\chi^{2}}(D_{e})^{2}=\sum_{e=1}^{\chi^{2}}(D^{\prime}_{e})^{2}=\sum_{e=1}^{\chi^{2}}(D^{\prime\prime}_{e})^{2}=1$. Since the spectra of the singular values, $\{D_{e}\}$, $\{D^{\prime}_{e}\}$, and $\{D^{\prime\prime}_{e}\}$, are different from each other, the EEs obtained from those singular values,
@@
are also different. Therefore, we can select the connection of the isometries with the smallest EE as the optimal one to update the local structure of TTN. Simultaneously, we can update the isometries in the central area by replacing them with the tensors obtained from the adopted singular-value decomposition, where we truncate the dimension of the center bond $e$ by keeping only the bases with the $\chi$-largest singular values.[32] We call the above procedures to update the local network structure and the isometries of the TTN [from Fig. 3(a) to Fig. 3(c)-(e)] a step. When this step is completed, we shift the central area by one isometry according to an appropriate sweeping rule of the network[16] and continue the computation. We iterate the steps by sweeping the entire network to obtain the optimal TTN. The algorithm is summarized in Table. 1.
@@
While both criteria using the EE of the center bond [Eq. (3)] or the truncated rDM weight [Eq. (4)] should work in searching for a good TTN structure, it is expected that using the latter is advantageous to improve the computational accuracy as much as possible with a fixed $\chi$ since the truncated rDM weight is directly related to the fidelity of the variational wave function in TTN with that of the true ground state. On the other hand, using the bond EE directly observes the entanglement and, therefore, is expected to be suitable to grasp the entanglement geometry of the ground state. It also bears noting that the algorithm using the truncated rDM weight should be strongly dependent on $\chi$.
@@
In the algorithm described above, one can improve only the isometries in the TTN with leaving the TTN structure fixed by always choosing the local connection $(ab|cd)$ for the singular-value decomposition Eq. (2a) at the process 5 in Table 1. Therefore, in practice, we can easily switch the calculations with and without the structural optimization of TTN. In our calculation discussed in the subsequent sections, we first perform the calculation with structural optimization at a certain $\chi=\chi_{\rm opt}$. Then, we continue the calculation without structural optimization with increasing $\chi$ to further improve the isometries in the TTN having the optimized structure.
In the singular-value decomposition of the wave function $\Psi(abcd)$, the number of non-zero singular values (the rank of the wave-function matrix in the decomposition considered) may be smaller than $\chi$. In such a case, only the bases corresponding to the non-zero singular values are adopted as the new bases for the center bond. In the practical calculation, bases with singular values smaller than a certain threshold (set to be $10^{-6}$ in our calculation) are discarded since their contribution to the wave function is negligibly small. In this case, which often occurs for the bonds close to the boundary of the TTN, the dimension of the bond is smaller than the upper bound $\chi$. Furthermore, in our calculation, all the bases whose singular values are degenerate are kept or discarded altogether to preserve symmetries of the target state. This measure may also result in an auxiliary bond with a dimension smaller than $\chi$.
In the calculation of the algorithm, one must be careful about the case that the dimensions of all three legs of an isometry become unity, which traps the calculation. Suppose that the dimensions of the bonds $a$, $b$, and $e$ are all unity in the situation shown in Fig. 3 (a). Then, the singular value of the connection with no structural change, shown in Fig. 3 (c), are $D_{1}=1$, $D_{e}=0$ $(e\geq 2)$, and the EE and truncated rDM weight are zero, which should be minimum among the three possible connections. As a result, the connection of Fig. 3 (c) will be selected, and the dimensions of the bonds $a$, $b$, and $e$ remain unity. (If the stochastic selection is adopted, there is a possibility that the other structure will be selected, but the original structure is most likely to be selected even in this case.) In such a way, the calculation can be stuck in the TTN containing the isometry whose three legs represent only one basis. In our calculation discussed in Sec. III, the problem occurred in some cases of the calculations of the Richardson model. We avoided the problem by taking an exceptional treatment in the preparation of the initial TTN. See Sec. III.3 and Appendix A for the details.

### v2

@@ Abstract
Numerical methods based on tensor networks have been extensively explored in the research of quantum many-body systems in recent years. It has been recognized that the ability of tensor networks to describe a quantum many-body state crucially depends on the spatial structure of the network. In the previous work [Hikihara et al., Phys. Rev. Res. 5, 013031 (2023)], we proposed an algorithm based on tree tensor networks (TTNs) that automatically optimizes the structure of TTN according to the spatial profile of entanglement in the state of interest. In this paper, we apply the algorithm to the random XY-exchange model under random magnetic fields and the Richardson model in order to analyze how the performance of the algorithm depends on the detailed updating schemes of the structural optimization. We then find that for the random XY model, on the one hand, the algorithm achieves improved accuracy, and the stochastic algorithm, which selects the local network structure probabilistically, is notably effective. For the Richardson model, on the other hand, the resulting numerical accuracy subtly depends on the initial TTN and the updating schemes. In particular, the algorithm without the stochastic updating scheme certainly improves the accuracy, while the one with the stochastic updates results in poor accuracy due to the effect of randomizing the network structure at the early stage of the calculation. These results indicate that the algorithm successfully improves the accuracy of the numerical calculations for quantum many-body states, while it is essential to appropriately choose the updating scheme as well as the initial TTN structure, depending on the systems treated.
@@ I Introduction
Since the development of the density-matrix renormalization-group (DMRG) method[1, 2], the tensor-network approach has made rapid progress in a variety of fields[3, 4, 5]. While tensor networks have been studied in statistical mechanics for many years[6], the introduction of quantum-information concepts has advanced the understanding of their properties significantly, and it has been recognized that tensor networks can represent the low-energy states of quantum many-body systems efficiently. Several algorithms based on various tensor-network states, including matrix-product state (MPS)[7, 8], tensor-product state (TPS)[9, 10] or projected entangled-pair state (PEPS)[11, 12], and the state of multiscale entanglement renormalization ansatz (MERA)[13, 14], have been developed.
@@
In this paper, we explore a numerical algorithm based on the tree-tensor networks (TTNs), which are tensor networks without loops. We particularly focus on binary trees, TTNs composed of three-leg tensors, although the extension to $m$-ary trees is straightforward. An essential property of the TTN is that cutting an arbitrary bond in the network separates the TTN into two subsystems. In other words, those two subsystems are connected only by the single bond. This property of TTN leads to the following guideline for reducing the loss of accuracy due to the truncation of the Hilbert space.[15, 16, 17] Since every bond in a TTN independently bridges the corresponding two subsystems, each bond must be able to express the entanglement between the subsystems in order to represent the target quantum state accurately. However, due to the upper bound of the bond dimension, which we denote by $\chi$ in the following, the entanglement entropy (EE) that a bond can carry is upper bounded to a maximum of $\ln\chi$. If the EE associated with each bond exceeds the upper limit, the excess entanglement is omitted in the approximated TTN and results in a loss of accuracy. Here, recall that the number of bonds in a TTN, which is 2N32N-3 where $N$ is the number of quantum spins in the system, is much smaller than the number of possible bipartitions of the system, $2^{N-1}-1$. Therefore, by using the network structure that makes the EEs on the (2N3)(2N-3) bonds in the TTN as small as possible, one can minimize the loss of accuracy in approximating the target quantum state by the TTN with a finite $\chi$. We call this guideline the least-EE principle.
@@
The least-EE principle indicates that using a TTN with the optimal structure is crucially important for improving the accuracy of the TTN approach. However, finding the optimal TTN structure is highly nontrivial and challenging. Several efforts have been devoted to the problem so far. There have been attempts to optimize the order of sites within the matrix-product network (MPN), which is employed in DMRG, based on various strategies.[18, 19, 20, 21, 22] A method to optimize the structure of TTN using the cut-off bond dimension as a minimization function has also been proposed.[23] For systems with randomness, the tensor-network strong-disorder renormalization group (tSDRG) method, in which a TTN suitable for a random system is constructed based on the idea of the strong-disorder renormalization group,[24, 25] has been developed[26, 27] and successfully applied.[28, 29, 15] For small systems for which an exact wavefunction is obtained, a method to extract the optimal TTN structure from the wavefunction according to the least-EE principle has also been proposed.[17]
@@
In the previous study[16], we proposed an algorithm to search for the optimal TTN structure by iterating the local reconstruction of the network based on the least-EE principle. We applied the algorithm to some toy models for which the optimal TTN structure was deduced from the perturbative renormalization analysis[16, 30]. We then demonstrated that the algorithm succeeded in obtaining the expected optimal structure. In this paper, we have investigated the performance of several types of TTN structural optimization calculations, which are specified by the initial TTN structure, the scheme to select the local TTN structure in the optimization process, and the bond dimension at which the structural optimization is executed. We have performed the calculations for the random XY-exchange model under random magnetic fields and the Richardson model. We have then demonstrated that for the random XY model, the structural optimization algorithm can reduce the variational energy significantly. In particular, a stochastic scheme, which selects the local TTN structure probabilistically, can achieve the lowest variational energy regardless of the initial TTN structure. For the Richardson model, on the other hand, we have found that while the structural optimization algorithm can yield a good TTN structure with a low variational energy, the calculations are sometimes trapped by a TTN structure unsuitable for representing the lowest energy state and fail to reach the optimal TTN structure. These results for the two models reveal that while the TTN structural optimization algorithm is basically effective in improving the accuracy of the calculation, the performance of each type of calculation depends on the model. The results thus indicate the importance of appropriately choosing the updating scheme of the network and the initial TTN structure according to the model to be studied.
@@
The plan of this paper is as follows. In Sec. II, we review the structural optimization algorithm and make several comments. Section III.1 describes the types of optimization calculations tested in the present study and the quantities calculated. Sections III.2 and III.3 respectively present the numerical results for the random XY-exchange model under random magnetic fields and the Richardson model. The performances of each type of calculation for the models are discussed there. Section IV is devoted to a summary and concluding remarks. In Appendix A, we explain how to prepare the initial TTN. The details of the actual calculation, including the number of sweeps, are presented in Appendix B. Some results of the calculations with a stochastic optimization scheme are presented in Appendix C. Appendix D describes the exact calculation of the lowest energy for the Richardson model.
@@ II Algorithm
Figure 1: Examples of TTN. (a) Matrix-Product Network (MPN), (b) Perfect-Binary Tree (PBT), (c) TTN of a generic form, and (d) the same TTN as that in (c). Open circles represent isometries and the solid diamond represents the singular-value matrix. Circles with colors represent bare spins. In (a), (b), and (c), the bare spins are arranged in the order of the site index. Note that within the category of MPN and PBT, there are several networks corresponding to different site orderings.
Figure 2: (a) Isometry $V_{a,b}^{c}$ with the legs $a$, $b$, and $c$. (b) Schematic picture of the orthonormal condition Eq. (1). Connected legs (bonds) represent a contraction of their degrees of freedom. The double arrow on the right-hand side of (b) represents a Kronecker delta $\delta_{cc^{\prime}}$.
@@
where $\bar{V}_{a,b}^{c}$ is a complex conjugate of $V_{a,b}^{c}$, and call the tensor the isometry. We assign a direction to each bond in the TTN so that the bases of the bond outgoing from each tensor are orthonormal; Namely, in Eq. (1), the index $c$ ($c^{\prime}$) corresponds to the outgoing bond, while the indices $a$ and $b$ correspond to the incoming bonds. All bonds in the TTN are directed from the boundary, where the bare spin degrees of freedom are, towards the “center” of the TTN, where the diagonal matrix of singular values is located. Such a representation format of the TTN is called the mixed canonical form[31]. Adopting the mixed canonical form offers strong advantages, such as guaranteeing the orthonormality of the basis of the truncated Hilbert space expanded by incoming bonds and making the contraction calculations of isometries extremely simple, as shown in Fig. 2 (b). We note that the center position of the canonical form can be taken at an arbitrary bond and can also be moved by the fusing and re-decomposing process of tensors[31].
Figure 3: (a) Central area at the beginning of a step. (b) Effective ground-state wave function $\Psi(abcd)$ obtained by the diagonalization of the effective Hamiltonian $\tilde{\mathcal{H}}$. (c), (d), and (e) show the possible local connections of the singular-value decompositions Eqs. (2a), (2b), and (2c), respectively. Open circles represent isometries and a solid diamond denotes the singular-value matrix at the canonical center.
@@
2. Construct the effective Hamiltonian $\tilde{\mathcal{H}}$ of the whole system on the orthonormal basis of the central area.
@@
4. Perform the singular-value decompositions of $\Psi$ for the three possible local connections in Figs. 3 (c)-(e).
@@ 5. Choose the optimal local connection according to the adopted updating scheme of the network discribed
9. Iterate the sweep until the TTN structure and quantities of interest (e.g., the ground-state energy) converge.
@@
Throughout this paper, we distinguish the isometries by the letters for the bond index, if necessary. [Hence, for example, $V^{e}_{ab}$ and $V^{e}_{cd}$ in Eq. (2a) are different isometries.] We also adopt the notation that the singular values are arranged in descending order, $D_{1}\geq D_{2}\geq\cdots\geq D_{\chi^{2}}\geq 0$ (the same for $\{D^{\prime}_{e}\}$ and $\{D^{\prime\prime}_{e}\}$) and suppose that the ground-state wave function $\Psi(abcd)$ is normalized so that $\sum_{e=1}^{\chi^{2}}(D_{e})^{2}=\sum_{e=1}^{\chi^{2}}(D^{\prime}_{e})^{2}=\sum_{e=1}^{\chi^{2}}(D^{\prime\prime}_{e})^{2}=1$. Since the spectra of the singular values, $\{D_{e}\}$, $\{D^{\prime}_{e}\}$, and $\{D^{\prime\prime}_{e}\}$, are different from each other, the EEs obtained from those singular values,
@@
are also different. Therefore, we can select the connection of the isometries with the smallest EE as the optimal one to update the local structure of TTN. Simultaneously, we can update the isometries in the central area by replacing them with the tensors obtained from the adopted singular-value decomposition, where we truncate the dimension of the center bond $e$ by keeping only the bases with the $\chi$-largest singular values.[32] We refer to the above procedures to update the local network structure and the isometries of the TTN [from Fig. 3(a) to Fig. 3(c)-(e)] as a “step”. When this step is completed, we shift the central area by one isometry according to an appropriate sweeping rule of the network[16] and continue the computation. We iterate the steps by sweeping the entire network to obtain the optimal TTN. The algorithm is summarized in Table. 1.
@@
It is expected that using Eq. (4) is advantageous to improve the computational accuracy as much as possible with a fixed $\chi$ since the truncated rDM weight is directly related to the fidelity of the variational wave function in TTN with the true ground-state wave function. On the other hand, using the bond EE [Eq. (3)] observes the entanglement and, therefore, is expected to be suitable to grasp the entanglement geometry of the ground state. It also bears noting that the algorithm using the truncated rDM weight should be strongly dependent on $\chi$.
@@
In the algorithm described above, one can improve only the isometries in the TTN while leaving the network structure fixed, by always choosing the local connection $(ab|cd)$ for the singular-value decomposition Eq. (2a) at process 5 in Table 1. Therefore, in practice, we can easily switch the calculations with and without the structural optimization of TTN. In our calculation discussed in the subsequent sections, we first perform the calculation with structural optimization at a certain $\chi=\chi_{\rm opt}$. Then, we continue the calculation without structural optimization by increasing $\chi$ to further improve the isometries in the TTN having the optimized structure.
In the singular-value decomposition of the wave function $\Psi(abcd)$, the number of non-zero singular values (the rank of the wave-function matrix in the decomposition considered) may be smaller than $\chi$. This situation often occurs for the bonds close to the boundary of the TTN. In such a case, only the bases corresponding to the non-zero singular values are adopted as the new bases for the center bond, and consequently, the dimension of the bond is smaller than the upper bound $\chi$. In the practical calculation, we discard the bases with singular values smaller than a certain threshold (set to be $10^{-6}$ in our calculation) since their contribution to the wave function is negligibly small. Furthermore, all the bases whose singular values are degenerate are kept or discarded altogether to preserve symmetries of the target state. These measures may also result in an auxiliary bond with a dimension smaller than $\chi$.
In the calculation, one must be careful about the case where the dimensions of all three legs of an isometry become unity, which traps the calculation. Suppose that the dimensions of the bonds $a$, $b$, and $e$ are all unity in the situation shown in Fig. 3 (a). Then, the singular values of the connection with no structural change, shown in Fig. 3 (c), are $D_{1}=1$, $D_{e}=0$ $(e\geq 2)$, and the EE and the truncated rDM weight are zero, which should be minimum among the three possible connections. As a result, the connection of Fig. 3 (c) will be selected, and the dimensions of the bonds $a$, $b$, and $e$ remain unity. (If the stochastic selection is adopted, there is a possibility that the other structure will be selected, but the original structure is most likely to be selected even in this case.) In such a way, the calculation can be stuck in the TTN containing the isometry whose three legs represent only one basis. In the present study, we encountered the problem in some cases of the calculations for the Richardson model. For such cases, we avoided the problem by taking an exceptional treatment in the preparation of the initial TTN. See Sec. III.3 and Appendix A for the details.